# A Complexity-Based Theory of Compositionality

### v1

@@ Appendix E Prequential coding
Instead of trying to estimate the terms on the LHS directly, we can estimate the RHS by finding the shortest program that jointly compresses both the dataset and the model, which we turns out to be easier through a compression algorithm called prequential coding illustrated in Figure E.1 and described below.
Figure E.1: Illustration of prequential coding, a method for estimating $K(D,\theta)=K(D|p_{\theta})+K(p_{\theta})$ using $p_{\theta}$’s learning algorithm $T$. a. Pseudocode of the prequential coding program that outputs both $D$ and $p_{\theta}$. The program jointly compresses $D$ and $p_{\theta}$ by incrementally training a model using $T$ on increasingly more data, each time efficiently encoding the next datapoint using the model obtained from all previous ones. The primary sources contributing to total program length come from specifying each next datapoint $D_{i+1}$ in compressed form using the current model $p_{\theta_{i}}$, which takes $-\log_{2}p_{\theta_{i}}(D_{i+1})$ bits. b. A visual illustration of the number of bits needed to specify each next datapoint given the model that was trained on all previous ones. As the learner $T$ sees more data, it outputs models that assign a higher likelihood to new observations, and can thus better compress them. The total prequential code length $L_{preq}(D;T)$ is given by the area under the curve. The area underneath the curve’s last point is equal to the number of bits needed to encode the entire dataset given the final model, $K(D|p_{\theta})$. Since $L_{preq}(D;T)=K(D|p_{\theta})+K(p_{\theta})$, the area above the curve’s last point is equal to $K(p_{\theta})$. Prequential coding formalizes the intuition that simple models generalize better, thus quickly decreasing their prediction error for the next datapoint.
@@
The total number of bits that it takes to jointly compress $D$ and $p_{\theta}$ using prequential coding is the sum of how many bits it takes to specify each next datapoint using a model that was trained on all previous ones. Visually, it is the area under the prequential coding curve shown in Figure E.1b. We can call the total length of this compression program the prequential code length $L_{preq}(D;T)$ (Blier and Ollivier,, 2018):
@@
Assuming that the model’s error decreases monotonically with the size of the training dataset, $K(D|p_{\theta})$ is equal to the area under the lowest point of the prequential coding curve in Figure E.1b. The area above this point is therefore the complexity of the model $K(p_{\theta})$. This relates Kolmogorov complexity to intuitions about generalization in ML: the simpler a model is, the quicker it generalizes from limited amounts of training data.
Appendix F Synthetic representations — experimental details
F.1 Lookup table representations
@@
F.2 Context-free grammar representations
@@
Appendix G Emergent languages — experimental details
@@
We estimated $K(Z|W^{L})$ for each language using prequential coding (see Appendix E). The model architecture used for prequential coding was an MLP with $2$ hidden layers of size $256$. Each word in $W^{L}$ embedded into a $64$-dimensional vector, and these concatenated embeddings were the input to the MLP. The MLP output logits over object values for each attribute. To estimate prequential code lengths more efficiently and avoid having to retrain the model $N$ times (where $N$ is the dataset size), we incremented the size of the dataset by chunks of size $50$ at a time. We used the Adam optimizer with a learning rate of $1\times 10^{-3}$ to train the model at each iteration of prequential coding. We reserved $400$ datapoints for a separate validation set that was used for early stopping at each iteration of prequential coding.
Appendix H Natural languages — experimental details
@@
We obtained English sentences from captions that were used to describe images in the Common Objects in Context (COCO) dataset (COCO,, 2024), downloaded from Hugging Face. The reason for using a dataset of image captions was that we expected these captions to use common words and simple sentence structures, given their grounding in visual stimuli. For each image, the dataset contained two independent captions, and we kept only the first. This gave us a total of $414,010$ English sentences. We then translated each sentence to French, Spanish, German, and Japanese using a large open-source language model with 3.3 billion parameters (Costa-jussà et al.,, 2022). We visually inspected several of the French, German, and Japanese sentences (no authors spoke Spanish) to make sure the translations were reasonable, and we found them to be of high quality. These sentences constituted the $W^{L}$’s for our experiments. We obtained proxies for the “meanings” $Z$ of these sentences by passing them through a large, pretrained, multilingual sentence embedding model that output a fixed-size vector for each sentence (Reimers and Gurevych,, 2020). Both the translation model and the sentence embedding model were obtained from Hugging Face.
@@
Estimating the compositionalities of these different language systems $C^{L}(Z)$ requires estimates of the numerator $K(Z)$ and denominator $K(Z|W^{L})$. While we did not estimate $K(Z)$, we assumed that it was approximately equal among languages. This is a common assumption in linguistics, where languages appear to be equivalent in their expressive power to express ideas, refer to objects, etc. Fixing the numerator $K(Z)$ to some (unknown) constant shared among languages allowed us to assess their relative compositionalities by estimating only the denominator $K(Z|W^{L})$. We estimated $K(Z|W^{L})$ for each language using prequential coding (see Appendix E).

### v3

@@ Appendix G Prequential coding
Instead of trying to estimate the terms on the LHS directly, we can estimate the RHS by finding the shortest program that jointly compresses both the dataset and the model, which we turns out to be easier through a compression algorithm called prequential coding illustrated in Figure G.1 and described below.
Figure G.1: Illustration of prequential coding, a method for estimating $K(D,\theta)=K(D|p_{\theta})+K(p_{\theta})$ using $p_{\theta}$’s learning algorithm $T$. a. Pseudocode of the prequential coding program that outputs both $D$ and $p_{\theta}$. The program jointly compresses $D$ and $p_{\theta}$ by incrementally training a model using $T$ on increasingly more data, each time efficiently encoding the next datapoint using the model obtained from all previous ones. The primary sources contributing to total program length come from specifying each next datapoint $D_{i+1}$ in compressed form using the current model $p_{\theta_{i}}$, which takes $-\log_{2}p_{\theta_{i}}(D_{i+1})$ bits. b. A visual illustration of the number of bits needed to specify each next datapoint given the model that was trained on all previous ones. As the learner $T$ sees more data, it outputs models that assign a higher likelihood to new observations, and can thus better compress them. The total prequential code length $L_{preq}(D;T)$ is given by the area under the curve. The area underneath the curve’s last point is equal to the number of bits needed to encode the entire dataset given the final model, $K(D|p_{\theta})$. Since $L_{preq}(D;T)=K(D|p_{\theta})+K(p_{\theta})$, the area above the curve’s last point is equal to $K(p_{\theta})$. Prequential coding formalizes the intuition that simple models generalize better, thus quickly decreasing their prediction error for the next datapoint.
@@
The total number of bits that it takes to jointly compress $D$ and $p_{\theta}$ using prequential coding is the sum of how many bits it takes to specify each next datapoint using a model that was trained on all previous ones. Visually, it is the area under the prequential coding curve shown in Figure G.1b. We can call the total length of this compression program the prequential code length $L_{preq}(D;T)$ (Blier and Ollivier,, 2018):
@@
Assuming that the model’s error decreases monotonically with the size of the training dataset, $K(D|p_{\theta})$ is equal to the area under the lowest point of the prequential coding curve in Figure G.1b. The area above this point is therefore the complexity of the model $K(p_{\theta})$. This relates Kolmogorov complexity to intuitions about generalization in ML: the simpler a model is, the quicker it generalizes from limited amounts of training data.
Appendix H Synthetic representations — experimental details
H.1 Lookup table representations
@@
H.2 Context-free grammar representations
@@
Appendix I Emergent languages — experimental details
@@
We estimated $K(Z|W^{L})$ for each language using prequential coding (see Appendix G). The model architecture used for prequential coding was an MLP with $2$ hidden layers of size $256$. Each word in $W^{L}$ embedded into a $64$-dimensional vector, and these concatenated embeddings were the input to the MLP. The MLP output logits over object values for each attribute. To estimate prequential code lengths more efficiently and avoid having to retrain the model $N$ times (where $N$ is the dataset size), we incremented the size of the dataset by chunks of size $50$ at a time. We used the Adam optimizer with a learning rate of $1\times 10^{-3}$ to train the model at each iteration of prequential coding. We reserved $400$ datapoints for a separate validation set that was used for early stopping at each iteration of prequential coding.
Appendix J Natural languages — experimental details
@@
We obtained English sentences from captions that were used to describe images in the Common Objects in Context (COCO) dataset (COCO,, 2024), downloaded from Hugging Face. The reason for using a dataset of image captions was that we expected these captions to use common words and simple sentence structures, given their grounding in visual stimuli. For each image, the dataset contained two independent captions, and we kept only the first. This gave us a total of $414,010$ English sentences. We then translated each sentence to French, Spanish, German, and Japanese using a large open-source language model with 3.3 billion parameters (Costa-jussà et al.,, 2022). We visually inspected several of the French, German, and Japanese sentences (no authors spoke Spanish) to make sure the translations were reasonable, and we found them to be of high quality. These sentences constituted the $W^{L}$’s for our experiments. We obtained proxies for the “meanings” $Z$ of these sentences by passing them through a large (278 million parameter), pretrained, multilingual sentence embedding model that output a fixed-size vector for each sentence (Reimers and Gurevych,, 2020). Both the translation model and the sentence embedding model were obtained from Hugging Face.
@@
Estimating the compositionalities of these different language systems $C^{L}(Z)$ requires estimates of the numerator $K(Z)$ and denominator $K(Z|W^{L})$. While we did not estimate $K(Z)$, we assumed that it was approximately equal among languages. This is a common assumption in linguistics, where languages appear to be equivalent in their expressive power to express ideas, refer to objects, etc. Fixing the numerator $K(Z)$ to some (unknown) constant shared among languages allowed us to assess their relative compositionalities by estimating only the denominator $K(Z|W^{L})$. We estimated $K(Z|W^{L})$ for each language using prequential coding (see Appendix G).